five persons A,B,C,D,E are pulling the cart of mass 100 kg on a smooth surface and cart is moving with acceleration of 3m/sq in east direction. when person A stops pulling it moves with acceleration of 1m/sq in west direction. When person B stops pulling, it moves with acceleration of 24m/sq in north direction. The magnitude of acceleration of the cart when only A and B pulls the cart keeping their direction same as the old direction is:

Grade:12

2 Answers

bharat bajaj IIT Delhi
9 years ago
F_{a} + F_{b} + F_{c} + F_{d} + F_{e} = 100. 3 i…. Eq. 1
where i is unit vector in east direction.
F_{b} + F_{c} + F_{d} + F_{e} = 100.-1i…. Eq 2
F_{a} + F_{c} + F_{d} + F_{e} = 100.24j… Eq 3
From Equation 1 and 2, we get
F_{a} = 100.4i
From equation 1 and 3 , we get
F_{b} = 100.(3i - 24j)
Hence, if only A and B are pulling,
F_{a} + F_{b} = 100.(7i - 24j)
Acceleration will be : 25 m/sq
Thanks & Regards
Bharat Bajaj
askiitian faculty
IIT Delhi
Tharun
5 years ago
 
F_{a} + F_{b} + F_{c} + F_{d} + F_{e} = 100. 3 i…. Eq. 1
where i is unit vector in east direction.
F_{b} + F_{c} + F_{d} + F_{e} = 100.-1i…. Eq 2
 
F_{a} + F_{c} + F_{d} + F_{e} = 100.24j… Eq 3
 
From Equation 1 and 2, we get
F_{a} = 100.4i
 
From equation 1 and 3 , we get
F_{b} = 100.(3i - 24j)
Hence, if only A and B are pulling,
F_{a} + F_{b} = 100.(7i - 24j)
Acceleration will be : 25 m/sq
Thanks & Regards
Bharat Bajaj
askiitian faculty
IIT Delh